OpenStudy (anonymous):

find the inverse of this function f(x)= -2x-5

OpenStudy (anonymous):

@gomathi

OpenStudy (anonymous):

this one says, multiply by minus two, subtract five inverse will say add five, divide by minus two

OpenStudy (anonymous):

just like if you were going to solve \[-2x-5=7\] you would add 5 and get \[-2x=12\] divide by -2 and get \[x=-6\]

OpenStudy (anonymous):

or you can do this, write \[y=-2x-5\] switch \(x\) and \(y\) and write \[x=-2y-5\] solve for \(y\) \[x+5=-2y\] \[\frac{x+5}{-2}=y\] and that is your inverse

OpenStudy (anonymous):

okay i understand the last one you put where you switch x and y. thanks!

OpenStudy (anonymous):

f=-2x-5 f+5=-2x -(f+5)/2=x inv 0f f=-(x+5)/2